OpenStudy (anonymous):

Determine which equation below has a solution of x = 2. A. 5x + 8 – 12x = –6x + 2x – 3 B. 6x + 4 – 12x = –10x + 18x – 3 C. –4x – 6 + x = –2x + x – 10 D. –3x – 8 + x = –4x + x – 10

OpenStudy (anonymous):

i mean C

OpenStudy (mathlegend):

Yes, off the top of my head I believe you are correct. How did you arrive at your answer? Did you do a plug in... or actually solve?

OpenStudy (anonymous):

A. 5(2)+8-12(2)= -6 -6(2)+2(2)-3= -11 FALSE B. 6(2)+4-12(2)= -8 -10(2)+18(2)-3= 13 FALSE C. –4(2) – 6 + 2 = -12 –2x + x – 10= -12 TRUE D. –3(2) – 8 + 2 = -12 –4x + x – 10= -16 FALSE So it's C!

OpenStudy (mathlegend):

As you can see @waterpologoalie chose to solve each one by simply plugging in the value of 2. Which is awesome or you could solve for x algebraically.

OpenStudy (anonymous):

i did what @waterpologoalie did

OpenStudy (anonymous):

sorry my computer freaked out and shut off

OpenStudy (anonymous):

@MathLegend is right, you could also solve for x first!